Two planets P and Q of the same mass.The radius of P is twice the radius of Q.vP and vQ are escape velocities from planet P and Q respectively.Find their escape velocities?


1
Expert's answer
2021-01-30T16:48:39-0500

According to the law of conservation of energy


"\frac{mv_Q^2}{2}-G\frac{Mm}{R}=0\to v_Q=\sqrt{2G\frac{M}{R}}"


"\frac{mv_P^2}{2}-G\frac{Mm}{2R}=0\to v_P=\sqrt{G\frac{M}{R}}"


"v_Q=\sqrt{2}\cdot v_P"